# Hawking–Page transition on a spin chain

### v1

@@ B.1 Improving the estimate of the BH entropy
Comparing (33) with the generalized spin chain matrix model (40), we observe that the entropy $S$, with irrelevant interactions [^1] In this letter, the word ‘irrelevant’ is always used in its precise RG flow sense. cut off at some order $K$, and the generalized spin chain with up to $K$ neighbour interactions, satisfy the analogous relation as for the simplified entropy $\hat{S}$ and the XX chain. Once again, the spin chain coupling constants $\tilde{J}_{n}$ are related to the Hubbard–Stratonovich fields $\sigma_{n}$ through $N\sigma_{n}=\tilde{J}_{n}/\tilde{T}\equiv J_{n}$, and in the generalized setting one is instructed to take a Gaussian average over all the spin-spin couplings $J_{n}$.
@@
B.2 Fermionic point of view
@@
$\mathcal{G}_{N}^{\text{\rm ferm}}=$ $\int_{0}^{2\pi}\frac{\mathrm{d}k_{1}}{2\pi}e^{-\epsilon_{k_{1}}/\tilde{T}}\cdots\int_{-\pi}^{\pi}\frac{\mathrm{d}k_{N}}{2\pi}e^{-\epsilon_{k_{N}}/\tilde{T}}$ (50)
@@
$\mathcal{G}^{\text{\rm ferm}}_{N}=\oint\mathrm{d}U\leavevmode\nobreak\ \exp\left\{\frac{1}{\tilde{T}}\mathrm{Tr}\leavevmode\nobreak\ V_{\epsilon}\left(U\right)\right\}$ (51)
@@
B.3 Perturbative corrections to the BH entropy
@@
$\hat{S}_{\lambda}\sim\ln\left[\int_{-\infty}^{\infty}\frac{\mathrm{d}\mu}{\mu}\leavevmode\nobreak\ e^{-\frac{N^{2}}{4b}(\mu-a)^{2}}\leavevmode\nobreak\ \left.e^{\hat{S}}\right\rvert_{a\mapsto\mu}\right]$ (57)
@@
B.4 Polyakov loop
From the point of view of the gauge theory, the Hawking–Page transition is holographically dual to a deconfinement transition at $T=T_{\rm HP}$ Witten (1998b). Famously, a probe of deconfinement ’t Hooft (1978) is provided by the so-called Polyakov loop Polyakov (1978), which is a Wilson loop winding around the Euclidean time direction, which is the circle $\mathbb{S}^{1}$ of radius $1/T$ in this context. Let us denote $\mathcal{P}$ the Polyakov loop expectation value, and let $A$ be the $SU(N)$ or $U(N)$ gauge field. Then,
where $\langle\cdot\rangle_{\mathcal{N}=4}$ stands for the correlation function in the full-fledged $\mathcal{N}=4$ CFT. Wilson loops in general, and $\mathcal{P}$ in particular, are charged under the centre symmetry of $\mathcal{N}=4$ SYM theory. This is directly seen by the fact that (59) is rotated by a phase when the gauge field $A$ is shifted by a flat 1-form in the centre of the gauge group. This makes $\mathcal{P}$ an order parameter for the deconfinement transition in $\mathcal{N}=4$ SYM: it must vanish in the low temperature phase, in which the centre symmetry is preserved, and acquires a vev in the high temperature phase, in which deconfinement spontaneously breaks the centre symmetry.
$\mathcal{P}=\left\langle\mathrm{Tr}U\right\rangle$ (60)
[…]
$\exp\left(N\varepsilon\mathrm{Tr}U\right)$ (61)
and, denoting $\hat{S}(\varepsilon)$ the entropy modified in this way,
On the spin chain side, it was shown in Pérez-García et al. (2022) that matrix model averages such as (60) correspond to impurities in the preparation of the initial state $\lvert\psi_{0}\rangle$. Then, $\mathcal{P}$ corresponds to the ratio of amplitudes with and without impurity. More precisely, let
$\lvert\psi_{\times}\rangle=\lvert\underbrace{\downarrow\downarrow\dots\downarrow}_{N-1}\uparrow,\downarrow,\uparrow\dots\uparrow\rangle,$ (64)
and denote $\mathcal{G}_{N}^{\times}(J)$ (respectively $\hat{\mathcal{G}}_{N}^{\times}(J)$) the analogue of $\mathcal{G}_{N}(J)$ (resp. $\hat{\mathcal{G}}_{N}(J)$) constructed using $\lvert\psi_{\times}\rangle$ as initial state, instead of $\lvert\psi_{0}\rangle$. Explicitly,
exactly as in (26). The centre symmetry of $\mathcal{N}=4$ SYM does act on the spin chain with the defect, rotating the amplitude $\hat{\mathcal{G}}_{N}^{\times}$ by a phase. However, any physical implementation would automatically select a real saddle point for $\mathcal{P}$, since the hypothetical experiment would only be measuring the real probability (left invariant by the action of the center symmetry) and not the amplitude. Stated differently, any concrete implementation of (66) would have built in the analogue of the regulating procedure described in Aharony et al. (2004) to obtain $\mathcal{P}\neq 0$.
In summary, the expectation value of a Polyakov loop can be analysed on the spin chain by repeating the same procedure and testing how the results vary when a simple impurity is introduced into the initial state.
It is also noteworthy that the definition of $\mathcal{P}$ can be modified in both $\mathcal{N}=4$ SYM and the spin chain. On the field theory side, it is possible to examine loops in higher-dimensional representations of the gauge group, such as the rank-$p$ antisymmetric representation. On the spin chain side, a feasible approach is to design an initial state $\lvert\psi_{\times}\rangle$ where the final spin $\downarrow$ is shifted $p\geq 1$ locations away from the other $N-1$ spins $\downarrow$. Despite this alteration, Identity (66) remains valid.
Let us note that, even without any reference to the gauge theory, we predict the ratio (67) of the Loschmidt echoes with and without impurities to be an order parameter discerning between a phase with low disorder, $0<a\ll 1$, and a phase of large disorder, thus large entropy, $a\gg 1$. This is a novel claim for the coupling averaged Heisenberg XX chain and may be of independent interest. To quantitatively test this prediction with the current state of the art technologies seems an interesting avenue to pursue.
@@ B.5 Complex temperature
In a laboratory setup, one first fixes the temperature of the system at a value $\tilde{T}_{\operatorname{\mathbb{R}}}$ and then let the system evolve until time $t$, which depends on the phase $\varphi$ of $a$ in such a way that
@@ B.6 Determinant identities from spin chains
The above argument is easily generalized to chain with $K^{\text{th}}$-nearest neighbour interaction, and no other interaction. Namely, one considers a spin chain with hopping Hamiltonian [^3] To reduce clutter, we slightly change the notation with respect to §B.1.
@@
[…]
@@ D.1 Operator algebras from the spin chain
It is interesting to ask whether other characteristic properties of the Hawking–Page transition are automatically encoded in the spin chain, if we look beyond the Loschmidt echo. For instance, phase transitions as the one studied presently are accompanied by a drastic change in the large $N$ algebra of operators Gesteau and Santilli (2024). Due to its relevance for questions concerning the BH horizon and its interior Leutheusser and Liu (2023a, b), it would be extremely interesting to enlarge our correspondence to include this aspect.
@@ D.2 Identifying states
So far, our conclusions have been about the large $N$ behaviour of the models, be it the gravitational system, its holographic dual $\mathcal{N}=4$ SYM or the spin chain. Nevertheless, the mapping of the effective partition function onto the spin chain is exact and thus holds for finite $N$. One problem, however, is that the contributions discarded in going from (6) to (8) give finite contributions at finite $N$, so their corrections should be taken into account. This can be done by constructing a more involved spin chain, with interactions beyond nearest neighbours as explained in §B.1, which approximates the effect of accounting for higher operators in (6). The situation will remain tractable for $N$ finite but large, but will become untamed for small $N$.
@@
This letter lays the groundwork for exploring important quantum simulation questions in the future. The exploration will shed light on the nature of black holes.

### v2

@@ B.2 Improving the estimate of the BH entropy
Comparing (42) with the generalized spin chain matrix model (49), we observe that the entropy $S$, with irrelevant interactions [^2] In this paper, the word “irrelevant” is always used in its precise RG flow sense. cut off at some order $K$, and the generalized spin chain with up to $K$ neighbour interactions, satisfy the analogous relation as for the simplified entropy $\hat{S}$ and the XX chain. Once again, the spin chain coupling constants $\tilde{J}_{n}$ are related to the Hubbard–Stratonovich fields $\sigma_{n}$ through $N\sigma_{n}=\tilde{J}_{n}/\tilde{T}\equiv J_{n}$, and in the generalized setting one is instructed to take a Gaussian average over all the spin-spin couplings $J_{n}$.
@@
B.3 Fermionic point of view
@@
$\mathcal{G}_{N}^{\text{\rm ferm}}=$ $\int_{0}^{2\pi}\frac{\mathrm{d}k_{1}}{2\pi}e^{-\epsilon_{k_{1}}/\tilde{T}}\cdots\int_{0}^{2\pi}\frac{\mathrm{d}k_{N}}{2\pi}e^{-\epsilon_{k_{N}}/\tilde{T}}$ (59)
@@
$\mathcal{G}^{\text{\rm ferm}}_{N}=\oint\mathrm{d}U\leavevmode\nobreak\ \exp\left\{\frac{1}{\tilde{T}}\mathrm{Tr}\leavevmode\nobreak\ V_{\epsilon}\left(U\right)\right\},$ (60)
@@
B.4 Perturbative corrections to the BH entropy
@@
$\hat{S}_{\lambda}\sim\ln\left[\int_{-\infty}^{\infty}\frac{\mathrm{d}\mu}{\mu}\leavevmode\nobreak\ e^{-\frac{N^{2}}{4b}(\mu-a)^{2}}\leavevmode\nobreak\ \left.e^{\hat{S}}\right\rvert_{a\mapsto\mu}\right],$ (66)
@@ B.5 Complex temperature
In a laboratory setup, one first fixes the temperature of the system at a value $\tilde{T}_{\operatorname{\mathbb{R}}}$ and then lets the system evolve until time $t$, which depends on the phase $\varphi$ of $a$ in such a way that
@@ B.6 Determinant identities from spin chains
The above argument is easily generalized to chain with $K^{\text{th}}$-nearest neighbour interaction, and no other interaction. Namely, one considers a spin chain with hopping Hamiltonian [^3] To reduce clutter, we slightly change the notation with respect to §B.2, $(\tilde{J}_{K}/K,\tilde{T})\mapsto(J,1)$.
@@
We conclude this appendix on a more speculative note. So far, our discussion has focused on a 1d quantum simulation to reproduce the entropy of a holographic BH in AdS5. We now suggest and lay down two open problems, which aim at looking for deeper signatures of BH physics on the spin chain.
@@ D.1 Operator algebras from the spin chain
It is interesting to ask whether other characteristic properties of the Hawking–Page transition are automatically encoded in the spin chain, if we look beyond the Loschmidt echo. For instance, phase transitions as the one studied presently are accompanied by a drastic change in the large-$N$ algebra of operators Gesteau and Santilli (2024). Due to its relevance for questions concerning the BH horizon and its interior Leutheusser and Liu (2023a, b), it would be extremely interesting to enlarge our correspondence to include this aspect.
@@ D.2 Identifying states
So far, our conclusions have been about the large-$N$ behaviour of the models, be it the gravitational system, its holographic dual $\mathcal{N}=4$ SYM, or the spin chain. Nevertheless, the mapping of the effective partition function onto the spin chain is exact and thus holds for finite $N$. One problem, however, is that the contributions discarded in going from (6) to (8) give finite contributions at finite $N$, so their corrections should be taken into account. This can be done by constructing a more involved spin chain, with interactions beyond nearest neighbours as explained in §B.2, which approximates the effect of accounting for higher operators in (6). The situation will remain tractable for $N$ finite but large, but will become untamed for small $N$.
@@
This paper lays the groundwork for exploring important quantum simulation questions in the future. The exploration will shed light on the nature of black holes.